### v1

# Explicit Constructions of Sum-Rank Metric Codes from Quadratic Kummer Extensions

Abstract
This paper presents new constructions of sum-rank metric codes derived from algebraic function fields, while the existing results about such codes remain limited. A central challenge in this field involves the determination of parameters. We address this challenge through quadratic Kummer extensions, and propose two general constructions of $2\times 2$ sum-rank codes. The novel sum-rank metric code constructions derived from algebraic function fields demonstrate superior code lengths compared to conventional coding paradigms, particularly when contrasted with linearized Reed-Solomon codes under equivalent code parameter constraints. We also establish explicit parameters including dimensions and minimum distances of our codes. Finally, an illustrative example through elliptic function fields is constructed for validating the theoretical framework.
Index terms: Sum-rank metric codes, algebraic geometry codes, function fields, elliptic curves.
I Introduction
Algebraic function fields over finite fields provide a foundational framework for Hamming-metric code construction [15]. Unified methodological frameworks are constructed through exploitation of algebraic structures inherent in function fields, enabling simultaneous characterization of both minimum distance parameters and dual code invariants [41]. Specifically, a linear $[n,k]$ Hamming metric code $C$ over $\mathbb{F}_{q}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum Hamming distance
$d_{H}:=\min_{c_{1}\neq c_{2}}\{wt_{H}(c_{1}-c_{2})\mid c_{1},c_{2}\in C\},$
where $wt_{H}$ counts non-zero coordinate positions. Classical algebraic geometry (AG) codes are constructed by evaluating functions in Riemann-Roch spaces at rational places, enabling explicit determination of code parameters.
For a linear $[n,k,d_{H}]$ code $C$, the Singleton bound asserts that
$d_{H}\leq n-k+1.$
When the equality holds, this code is called a maximum distance separable (MDS) code. Reed-Solomon (RS) codes [18], derived from rational function fields $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ with $x$ transcendental over $\mathbb{F}_{q}$, represent prominent Hamming metric MDS codes. Their maximal length of $q+1$ over $\mathbb{F}_{q}$ coincides with the main conjecture of MDS codes.
Very recently, sum-rank metric codes have attracted significant attention due to applications in multishot network coding [22, 24, 31, 35], space-time coding [40], and distributed storage [23, 25]. For more applications and details, we refer the reader to [27]. The basic definitions are given as follows.
Definition I.1.
Given $s\geq 1$ and positive integers $n_{1},\ldots,n_{s},m_{1},\ldots,m_{s}$ with $n_{i}\leq m_{i}$ and $m_{1}\geq m_{2}\cdots\geq m_{s}$, denote
$\Pi:=\Pi_{q}(m_{1}\times n_{1}|\cdots|m_{s}\times n_{s})=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{m_{i}\times n_{i}}$
where $\mathbb{F}_{q}^{m_{i}\times n_{i}}$ denotes the space of $m_{i}\times n_{i}$ matrices over $\mathbb{F}_{q}$. Note that $\Pi$ forms an $\mathbb{F}_{q}$-linear space of dimension $\sum_{i=1}^{s}m_{i}n_{i}$. The sum-rank weight of
$X:=(X_{1},\ldots,X_{s})\in\Pi$
is defined as
$wt_{sr}(X):=\sum\limits_{i=1}^{s}rank(X_{i}).$
The sum-rank distance between $X,Y\in\Pi$ is
$d_{sr}(X,Y)=wt_{sr}(X-Y).$
The sum-rank weight and distance induce a metric on $\Pi$. We formally define a linear sum-rank metric code as follows.
Definition I.2.
A linear sum-rank metric code $C_{sr}$ is an $\mathbb{F}_{q}$-linear subspace of $\Pi$ with minimum sum-rank distance
$d_{sr}(C):=\min_{X\neq Y}\{wt_{sr}(X-Y)|X,Y\in\Pi\}.$
For notational conciseness, we adopt the convention of referring to sum-rank metric codes as sum-rank codes throughout this work. The sum-rank code framework constitutes a natural generalization of classical coding systems: (i) When configured with block uniformity parameters $n_{i}=m_{i}=1$ for all $1\leq i\leq s$, these codes precisely recover the Hamming metric codes of length $s$ through coordinate-wise symbol comparison; (ii) In the monolithic regime where $s=1$, the construction collapses to the original rank-metric codes introduced by Delsarte [12]. This multifunctional parametrization reveals the intrinsic capacity of the codes to interpolate between symbol-level and matrix-level error protection strategies.
The theoretical landscape of sum-rank codes is underpinned by three foundational pillars as established in prior research:
  1. 1.
    Structural Properties: Code duality mechanisms and weight enumerator analysis [1, 5]
  2. 2.
    Distance Bounds: Fundamental limitations including Singleton-type and Johnson-type bounds [37]
  3. 3.
    Decoding Thresholds: Characterization of list-decoding capacity regions [2]
The Singleton bound for sum-rank codes was provided in [5] and [25]. For the general case, let $j$ and $\delta$ be the unique integers satisfying $d-1=\sum_{i=1}^{j-1}n_{i}+\delta$ with $10\leq\delta\leq n_{j}-$ 1. Then
$|C|\leq q^{\sum_{i=j}^{s}m_{i}n_{i}-m_{j}\delta}.$
A code $C\subset\Pi$ attaining this bound is called maximum sum-rank distance (MSRD). For the special case $n_{1}=\cdots=n_{s}=n$ and $m_{1}=\cdots=m_{s}=m$, this bound is of the form
$|C|\leq q^{m(sn-d_{sr}+1)}.$
Throughout this paper, we focus on sum-rank metric codes with these uniform parameters.
Constructions of “good” sum-rank metric codes can be found in [1, 6, 7, 9, 23, 26, 28, 29, 30, 33, 34]. A standard technique constructs $C$ as a classical Hamming metric code of length $sn$ over $\mathbb{F}_{q^{m}}$ and then applies rank metric code constructions to each $n$-length block. This ensures that the dimension $\kappa$ of $C$ over $\mathbb{F}_{q^{m}}$ satisfies $d_{sr}+\kappa\leq sn+1$, matching the Singleton bound for the Hamming metric. Moreover, efficient algorithms over $\mathbb{F}_{q^{m}}$ can be employed for decoding sum-rank metric codes, see [3, 10, 16, 39].
One of the most prominent rank metric codes is the Gabidulin code, constructed using linearized polynomials over $\mathbb{F}_{q^{m}}$, see [13, 14, 19, 21, 32]. Among sum-rank codes, linearized Reed-Solomon codes [23] represent the most extensively investigated families to date, see [8, 17, 20, 33, 38]. Gabidulin codes and linearized RS codes are maximum rank distance (MRD) and MSRD respectively, serving as rank metric and sum-rank metric analogues of RS codes.
Motivations
While existing algebraic geometric constructions [9, 28] for sum-rank metric codes are non-trivial, most results are constructed using towers of function fields as illustrated in Figure 1. In [4], E. Berardini and X. Caruso established the first geometric construction of sum-rank metric codes. Their approach employs the Ore polynomial ring [36] over an extension field derived from the completion of $F$ at specified places and subsequently develops the theory of Riemann-Roch spaces over this ring. This framework enables them to introduce linearized algebraic geometry codes and presents explicit constructions of sum-rank metric codes with parameters consistent with classical algebraic geometry codes.
A principal motivation for constructing codes via algebraic function fields lies in their ability to provide numerous evaluation places bounded by the Hasse-Weil bound with higher genus. Consequently, codes with long lengths can be constructed over small finite fields. Classical algebraic geometry codes are known to asymptotically surpass the Gilbert-Varshamov (GV) bound [43]. Recent work [37] demonstrates that random linear sum-rank metric codes attain the asymptotic GV-like bound with high probability. This raises a natural question: whether sum-rank codes derived from algebraic geometry also possess parameters surpassing the GV bound.
[…]
Figure 1: Tower of function field extensions.
Our contributions
In contrast to the construction in [4], we propose an alternative method to construct sum-rank metric codes via function fields. The key tool of our construction is the local integral basis of a place in function field extensions. Specifically, let $\mathcal{O}_{P}$ be the valuation ring at a place $P$ of $F/\mathbb{F}_{q}$, and $F^{\prime}/F$ a finite separable function field extension. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ forms a vector space over $\mathcal{O}_{P}$. Consequently, a function in $F^{\prime}/\mathbb{F}^{\prime}_{q}$ that is regular at the places lying over $P$ corresponds to a matrix over $F$, where $\mathbb{F}^{\prime}_{q}$ is a finite extension of $\mathbb{F}_{q}$. When $F^{\prime}/F$ is Galois, we establish a linearized-polynomial-like map on $\mathcal{O}^{\prime}_{P}$ through the direct sum of Riemann-Roch spaces, yielding sum-rank metric codes via evaluation of the function matrix at rational places. While our construction overlaps with [4], it differs substantially by emphasizing the intrinsic structure of function field extensions. By sacrificing certain subspaces in the direct sum of Riemann-Roch spaces, we explicitly determine the parameters of our sum-rank codes.
Organization
The remainder of this paper is organized as follows. Section II introduces the foundational concepts necessary for our purpose. In Section III, we establish a general construction of sum-rank metric codes through function field extensions. Section IV presents the explicit sum-rank metric codes via quadratic Kummer extensions and introduces a new specialized construction, including a case of elliptic function fields. Section V concludes this paper.
II Preliminaries and basic concepts
This section reviews some basic definitions and notations, primarily following [41].
II-A Places and divisors
Let $q$ be a prime power. An algebraic function field $F/\mathbb{F}_{q}$ denotes a finite extension of $\mathbb{F}_{q}(x)$ with genus $g(F)$, where $x$ is transcendental over $\mathbb{F}_{q}$. A place $P$ is the maximal ideal of a valuation ring $\mathcal{O}_{P}$ in $F/\mathbb{F}_{q}$ with
$\deg(P)=[\mathcal{O}_{P}/P:\mathbb{F}_{q}].$
If $\deg(P)=1$, then $P$ is called rational. The prime element of $P$ is a function $t\in P$ such that $P=t\mathcal{O}_{p}$. Any non-zero $z\in F$ has a unique representation $z=t^{n}u$ with $u\in\mathcal{O}^{*}_{P}$. The associated discrete valuation
$v_{P}:F\to\mathbb{Z}\cup\{\infty\}$
satisfies $v_{P}(z)=n$ and $v_{P}(0)=\infty$. The strict triangle inequality of $v_{P}$ is characterized as follows:
Lemma II.1.
For any discrete valuation $v_{P}$,
$v_{P}(x+y)\geq\min\{v_{P}(x),v_{P}(y)\}$
holds for all $x,y\in F$. Moreover, if $v_{P}(x)\neq v_{P}(y)$, then $v_{P}(x+y)=\min\{v_{P}(x),v_{P}(y)\}$.
A divisor $G$ is defined as the formal sum of places
$G:=\sum\limits_{P}n_{P}P$
with $n_{P}\in\mathbb{Z}$ and $\deg(G)=\sum\limits_{P}n_{P}$. For any function $z\in F$,
$(z):=\sum\limits_{P}v_{P}(z)P$
is called the principal divisor. Given two divisors $G_{1},G_{2}$, we write $G_{1}\leq G_{2}$ if $n_{P}(G_{1})\leq n_{P}(G_{2})$ for all places. The associated Riemann-Roch space is defined as
$\mathcal{L}(G):=\{z\in F\mid(z)+G\geq 0\}\cup\{0\}.$
The space is a vector space over $\mathbb{F}_{q}$ with dimension given by the Riemann-Roch theorem:
Lemma II.2.
[41, Theorem 1.4.17] Let $\ell(G)$ denote the dimension of $\mathcal{L}(G)$ over $\mathbb{F}_{q}$. Then
$\ell(G)\geq\deg(G)-g(F)+1.$
Moreover, if $\deg(G)\geq 2g(F)-1$, then
$\ell(G)=\deg(G)-g(F)+1.$
II-B Extensions of function fields
A finite algebraic function field $F^{\prime}/\mathbb{F}_{q^{m}}$ is called an algebraic extension of $F/\mathbb{F}_{q}$ if $F^{\prime}/F$ is an algebraic field extension with $[F^{\prime}:F]<\infty$ and $m\geq 1$. A place $P^{\prime}$ of $F^{\prime}$ is said lying over $P$ if $P\subseteq P^{\prime}$, denoted $P^{\prime}|P$. The ramification index $e(P^{\prime}|P)$ is defined as the integer satisfying
$v_{P^{\prime}}(x)=e(P^{\prime}|P)\cdot v_{P}(x)$
for all $x\in F$, while the relative degree of $P^{\prime}$ over $P$ is given by
$f(P^{\prime}|P)=[\mathcal{O}_{P^{\prime}}/P^{\prime}:\mathcal{O}_{P}/P].$
For any place $P$ of $F$, there exist only finitely many places $P^{\prime}$ of $F^{\prime}$ with $P^{\prime}|P$. The conorm of $P$ is defined as
$\text{Con}_{F^{\prime}/F}(P):=\sum\limits_{P^{\prime}|P}e(P^{\prime}|P)\cdot P^{\prime}$
and
$\text{Con}_{F^{\prime}/F}\left(\sum\limits_{P}n_{P}P\right):=\sum n_{P}\cdot\text{Con}_{F^{\prime}/F}(P).$
Consequently, for any non-zero function $z\in F$, the principal divisor of $z$ in $F^{\prime}$ has the form
$(z)_{F^{\prime}}=\text{Con}_{F^{\prime}/F}((z)_{F}).$
The fundamental equality of places is stated as follows.
Proposition II.1.
[41, Theorem 3.3.11] Let $P$ be a place of $F$, and let $P_{1},\ldots,P_{n}$ be all places of $F^{\prime}$ lying over $P$. Denote $e_{i}=e(P_{i}|P)$ and $f_{i}=f(P_{i}|P)$. Then
$\sum\limits_{i=1}^{n}e_{i}f_{i}=[F^{\prime}:F].$
If $[F^{\prime}:F]=n$ and there exist exactly $n$ distinct places lying over $P$, then $P$ splits completely in $F^{\prime}/F$. In this case, it follows that $e_{i}=f_{i}=1$. When $e(P^{\prime}|P)>1$ for some $P^{\prime}$, the place $P$ is ramified in $F^{\prime}/F$. Moreover, if there exists a place $P^{\prime}$ with $e(P^{\prime}|P)=n$, then $P$ is totally ramified in $F^{\prime}/F$.
II-C Galois Extensions and Kummer Extensions
An extension $F^{\prime}/\mathbb{F}_{q^{m}}$ of a function field $F/\mathbb{F}_{q}$ is Galois if $F^{\prime}/F$ is a finite Galois extension. Let $\text{Gal}(F^{\prime}/F)$ be the Galois group of $F^{\prime}/F$. For any place $P$ of $F$, the Galois group acts transitively on the set of places lying over $P$.
Kummer extensions constitute a special class of Galois extensions. We present their definition and key properties.
Definition II.1.
Let $\mathbb{F}_{q}$ contain a primitive $n$-th root of unity $\zeta$. Suppose that $u\in F$ satisfies
$u\neq\omega^{d}$
for all $\omega\in F$ and $1<d|n$. Let $F^{\prime}=F(y)$ with $y^{n}=u$ be an extension of $F/\mathbb{F}_{q}$. Then $F^{\prime}$ is called a Kummer extension of $F$ with:
  • $\text{Gal}(F^{\prime}/F)=\langle\sigma\rangle$ where $\sigma(y)=\zeta y$;
  • For any place $P$ of $F$ and $P^{\prime}$ lying over $P$, $e(P^{\prime}|P)=\frac{n}{\gcd(n,v_{P}(u))}$;
  • If there exists a place $Q$ of $F$ with $\gcd(n,v_{Q}(u))=1$, then $\mathbb{F}_{q}$ remains the constant field of $F^{\prime}$, and
    $g(F^{\prime})=1+n(g+1)+\frac{1}{2}\sum\limits_{P}(n-\gcd(n,v_{P}(u)))\deg(P);$
  • For any place $P$ with $v_{P}(u))\geq 0$, the basis $\{1,y,\ldots,y^{n-1}\}$ is a local integral basis of $F^{\prime}/F$ for $P$.
III A General Setting of Sum-Rank Metric Codes
In this section, we present a general construction of sum-rank metric codes through function field extensions, which differs from the approach by E. Berardini and X. Caruso in [4].
III-A The Main Idea of Matrices over $\mathbb{F}_{q}$
Let $F/\mathbb{F}_{q}$ be a function field and $F^{\prime}/\mathbb{F}_{q^{m}}$ a Galois extension of $F/\mathbb{F}_{q}$ with $[F^{\prime}:F]=n$. Suppose that $P$ is a place of $F$ and $t\in F$ is a prime element of $P$. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ is given by
$\mathcal{O}^{\prime}_{P}=\bigcap\limits_{i=1}^{n}\mathcal{O}_{P_{i}}.$
An integral basis $\{z_{1},\ldots,z_{n}\}$ of $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is a basis of $F^{\prime}/F$ satisfying
$\mathcal{O}^{\prime}_{P}=\sum\limits_{i=1}^{n}\mathcal{O}_{P}z_{i}.$
The quotient $V:=\mathcal{O}^{\prime}_{P}/t\mathcal{O}^{\prime}_{P}$ forms a vector space over $\mathcal{O}_{P}/P$ with basis
$\{z_{1}+t\mathcal{O}^{\prime}_{P},\ldots,z_{n}+t\mathcal{O}^{\prime}_{P}\}.$
For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and place $P^{\prime}$ with $P^{\prime}|P$, observe that for $z\in\mathcal{O}^{\prime}_{P}$:
$v_{P^{\prime}}(\sigma(z))=v_{\sigma^{-1}(P^{\prime})}(z)\geq 0$
since $\sigma^{-1}(P^{\prime})|P$. This implies $\sigma(z)\in\mathcal{O}^{\prime}_{P}$, hence $\sigma(V)=V$. The set
$\{\sigma(z_{1})+t\mathcal{O}^{\prime}_{P},\ldots,\sigma(z_{n})+t\mathcal{O}^{\prime}_{P}\}$
forms another basis of $V$ over $\mathcal{O}_{P}/P$. Then $\sigma$ induces a matrix $\bar{A}_{\sigma}$ over $\mathcal{O}_{P}/P$ with respect to $\{z_{i}\}$.
Similarly, for $f\in\mathcal{O}^{\prime}_{P}$, consider the multiplication $\mu_{f}:\mathcal{O}^{\prime}_{P}\to\mathcal{O}^{\prime}_{P}$ defined by $\mu_{f}(z)=fz$. This induces an $\mathcal{O}_{P}/P$-linear map $\bar{\mu}_{f}:V\to V$ via
$\bar{\mu}_{f}(z+t\mathcal{O}^{\prime}_{P}):=fz+t\mathcal{O}^{\prime}_{P}.$
This induces a matrix $\bar{B}_{f}$ with entries in $\mathcal{O}_{P}/P$.
Remark III.1.
The matrix $A=(a_{ij})$ associated to $\sigma$ has entries in $\mathcal{O}_{P}$ due to the integral basis property. Under the reduction map
$\pi:\mathcal{O}_{P}\to\mathcal{O}_{P}/P,$
we obtain $\bar{A}=(\pi(a_{ij}))$. This explains the notation $\bar{A}_{\sigma}$ and $\bar{B}_{f}$.
When $P$ is rational (i.e., $\mathcal{O}_{P}/P=\mathbb{F}_{q}$), both $\sigma\in\text{Gal}l(F^{\prime}/F)$ and $f\in\mathcal{O}^{\prime}_{P}$ induce matrices over $\mathbb{F}_{q}$. Let $\text{Gal}(F^{\prime}/F)=\{\sigma_{1},\ldots,\sigma_{n}\}$, and consider functions $f_{1},\ldots,f_{n}\in\mathcal{O}^{\prime}_{P}$. We define a linear operator
$L:\mathcal{O}^{\prime}_{P}$ $\to\mathcal{O}^{\prime}_{P}$
$g$ $\mapsto\sum\limits_{i=1}^{n}f_{i}\sigma_{i}(g).$
This $F^{\prime}$-linear operator naturally induces a linear morphism on $F$. We further define a map $\varepsilon_{P}$ satisfying
$\varepsilon_{P}(L)=\sum\limits_{i=1}^{n}A_{i}B_{i}=(M_{ij})$
where $B_{i}$ and $A_{i}$ are matrices representing $\mu_{f_{i}}$ and $\sigma_{i}$ for $1\leq i\leq n$, respectively. Furthermore, let
$\bar{\varepsilon}_{P}(L)=\sum\limits_{i=1}^{n}\bar{A_{i}}\bar{B_{i}}=(\pi(M_{ij})).$
We establish the following lemma.
Lemma III.1.
  • (i)
    Let $\{z_{1},\ldots,z_{n}\}$ be an integral basis of $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$. Then the matrix representing $L$ with respect to this basis is $\varepsilon_{P}(L)$.
  • (ii)
    The determinant $\det(\varepsilon_{P}(L))$ belongs to $\mathcal{O}_{P}$.
  • (iii)
    The kernel ${\rm ker}(L)\neq\{0\}$ if and only if $\det(\varepsilon_{P}(L))=0$.
Proof.
(i) Let $g=g_{1}z_{1}+\cdots+g_{n}z_{n}\in\mathcal{O}^{\prime}_{P}$. For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and $f\in\mathcal{O}_{P}$, let $A_{\sigma}$ and $B_{f}$ denote the matrices defined above. Then
$\mu_{f}(g)=[g_{1},\ldots,g_{n}]B_{f}[z_{1},\ldots,z_{n}]^{T}$
and
$\sigma(g)=[g_{1},\ldots,g_{n}]A_{\sigma}[z_{1},\ldots,z_{n}]^{T}.$
Thus $f\sigma(g)=[g_{1},\ldots,g_{n}]B_{f}A_{\sigma}[z_{1},\ldots,z_{n}]^{T}$. The conclusion follows from the linearity of $\mu_{f}$ and $\sigma$.
(ii) The result immediately follows since all coefficients of $A_{\sigma_{i}}$ and $B_{f_{i}}$ are contained in $\mathcal{O}_{P}$.
(iii) Consider the equation
$\varepsilon_{P}(L)^{T}[x_{1},\ldots,x_{n}]^{T}=[0,\ldots,0].$
Any solution $[g_{1},\ldots,g_{n}]$ satisfies $L(g_{1}z_{1}+\cdots+g_{n}z_{n})=0$. Therefore ${\rm ker}(L)$ contains non-zero elements if and only if the system has non-trivial solutions, which occurs precisely when $\det(\varepsilon_{P}(L))=0$. ∎
III-B Sum-rank Codes in General Case
Now we present the construction of sum-rank metric codes. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$. Assume that the divisor $Q_{0}$ is totally ramified in $F^{\prime}/F$ with $P_{0}|Q_{0}$. Let $G=nQ_{0}$, $G^{\prime}=nP_{0}$ and $D=Q_{1}+\cdots+Q_{s}$. Define the evaluation map:
$\text{ev}_{D,1}:\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$ $\to\Pi$
$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
Since $f_{1},\ldots,f_{n}$ have poles only at $P_{0}$, they belong to $\mathcal{O}^{\prime}_{Q_{i}}$ for all $1\leq i\leq s$. Thus $\varepsilon_{Q_{i}}(L)$ are well-defined. Moreover, the function space $\mathcal{L}(G^{\prime})$ forms an $\mathbb{F}_{q}$-linear space and $L$ is $F$-linear. It follows that $\text{ev}_{D,1}$ is an $\mathbb{F}_{q}$-linear map. Thus the image of $\text{ev}_{D,1}$ is a subspace of $\Pi$. The code $C_{sr,1}(D,G)$ is defined as the image of $\text{ev}_{D,1}$.
Remark III.2.
The code $C_{sr,1}$ resembles the code $C(1,G,Q_{1},\ldots,Q_{s})$ defined in [4, Definition 4]. In their work, E. Berardini and X. Caruso constructed codes through the isomorphism chain
$D_{F^{\prime}_{i},x}\to D_{F^{\prime}_{i},1}\to\text{End}_{F_{i}}(F^{\prime}_{i})$
where $F_{i}$ denotes the completion of $F$ at $Q_{i}$, while $F^{\prime}_{i}$ represents the product of completions of $F^{\prime}$ at places lying over $Q_{i}$ for each $i=1,\ldots,n$. Our approach directly employs $L$ to express linearized-polynomial-like maps, corresponding to elements of $D_{F^{\prime},1}$. We subsequently utilize the integral closure $\mathcal{O}^{\prime}_{Q_{i}}$ to construct explicit matrices $\bar{\varepsilon}_{Q_{i}}(L)$, where $\varepsilon_{Q_{i}}(L)$ is induced by $L$ at $Q_{i}$.
Furthermore, while previous work proposed several hypotheses for code parameter determination, the ring $D_{F^{\prime}_{i},1}$ fails to satisfy all necessary conditions. In subsequent sections, we provide more explicit dimension and minimum distance bounds by considering subspaces of $\bigoplus_{i=1}^{n}\mathcal{L}(G^{\prime})$.
Proposition III.1.
The length $n^{2}s$ of $C_{sr,1}$ over $\mathbb{F}_{q}$ satisfies
$n^{2}s\leq n^{2}(q+2g(F)\sqrt{q}).$
Proof.
In the construction of $C_{sr,1}$, each rational place $Q\neq Q_{0}$ yields a well-defined $n\times n$ matrix $\bar{\varepsilon}_{P}(L)$ over $\mathbb{F}_{q}$. By the Hasse-Weil bound, the function field $F/\mathbb{F}_{q}$ contains at most $q+1+2g(F)\sqrt{q}$ rational places, implying $s\leq q+2g(F)\sqrt{q}$. ∎
IV Sum-Rank Metric Codes via Quadratic Extension of Algebraic Function Fields
IV-A Parameters of the General Construction
We first establish some necessary lemmas before deriving the dimension and minimum sum-rank distance for $C_{sr,1}$.
Lemma IV.1.
Suppose that $Q\neq P$ is a rational place of $F$ with $Q$ totally ramified in $F^{\prime}/F$. Let $Q^{\prime}$ be the place of $F^{\prime}$ lying over $Q$. For $f_{1},\ldots,f_{n}\in\mathcal{L}(kQ^{\prime})$, we have
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Furthermore, if there exists $f_{i}$ satisfying
$v_{Q^{\prime}}(f_{i})<v_{Q^{\prime}}(f_{j}),\forall j\neq i,$
then $\det(\varepsilon_{P}(L))\neq 0$.
Proof.
Since $f_{i}\in\mathcal{L}(kQ^{\prime})$, the coefficients of $B_{f_{i}}$ are contained in $\mathcal{L}(kQ^{\prime})$ for $i=1,\ldots,n$. The discrete valuation of $\det(\varepsilon_{P}(L))$ at $Q^{\prime}$ therefore exceeds the minimal valuation in its determinant expansion. Consequently, we have $\det(\varepsilon_{P}(L))\in\mathcal{L}(knQ^{\prime})$. Since $Q$ is totally ramified, the conorm of $Q$ is
$\text{Con}_{F^{\prime}/F}(Q)=nQ^{\prime}.$
By Lemma III.1 (ii) and the conorm properties, we have
$v_{Q}(\det(\varepsilon_{P}(L)))\leq k.$
Note that for any $Q^{\prime}$ of $F^{\prime}$, there exists only one $Q$ in $F$ with $Q^{\prime}|Q$, thus $\det(\varepsilon_{P}(L))$ only has poles at $Q$. Therefore $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ)$. For any non-zero $g\in\mathcal{O}^{\prime}_{P}$, the equality
$v_{Q^{\prime}}(\sigma_{i}(g))=v_{Q^{\prime}}(\sigma_{1}(g))$
holds for $i=1,\ldots,n$. W.l.o.g, assume that $v_{Q^{\prime}}(f_{1})<v_{Q^{\prime}}(f_{j})$ when $2\leq j\leq n$. This yields
$v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(f_{j}\sigma_{j}(g)).$
By the strict triangle inequality, we have
$v_{Q^{\prime}}\left(\sum\limits_{j=1}^{n}f_{j}\sigma_{j}(g)\right)=v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(\sigma_{1}(g)).$
Thus $L(g)\neq 0$ for $g\neq 0$. The conclusion follows from Lemma III.1 (iii). ∎
Lemma IV.2.
Let $P$ be a place of $F$ with a prime element $t$. If $\det(\varepsilon_{P}(L))\neq 0$ for $L$ associated tp $Q\neq P$ while $rank(\bar{\varepsilon}_{P}(L))<n$, then $\det(\varepsilon_{P}(L))\in P$.
Proof.
The condition $rank(\bar{\varepsilon}_{P}(L))<n$ means $\det(\bar{\varepsilon}_{P_{\alpha}}(L))=0$. Let $\pi_{P}$ denote the residue class map at $P$. For $\varepsilon_{P}(L)=(M_{ij})$, we have
$\bar{\varepsilon}_{P}(L)=(\pi_{P}(M_{ij}))=(M_{ij})(P).$
Therefore
$\det(\bar{\varepsilon}_{P}(L))=\det(\varepsilon_{P}(L))(P).$
Since $\det(\varepsilon_{P}(L))(P)=0$, it follows that $\det(\varepsilon_{P}(L))\in P$. ∎
Lemma IV.3.
Let $k,k_{1}$ be integers with $k>k_{1}\geq 2g(F)-1$, and $P$ a rational place of $F$. Define
$\mathcal{L}^{*}_{k_{1}}(kP):=\{f|-k\leq v_{P}(f)<-k_{1}\}\cup\{0\}.$
Then $\mathcal{L}^{*}_{k_{1}}(kP)$ forms a vector space over $\mathbb{F}_{q}$ of dimension $k-k_{1}$.
Proof.
By the Riemann-Roch theorem, we obtain that $\mathcal{L}(kP)$ and $\mathcal{L}(k_{1}P)$ are vector spaces over $\mathbb{F}_{q}$ with dimensions $\ell_{k}:=k-g(F)+1$ and $\ell_{k_{1}}:=k_{1}-g(F)+1$, respectively. Thus,
$\mathcal{L}^{*}_{k_{1}}(kP):=\mathcal{L}(kP)\setminus\mathcal{L}(k_{1}P)\cup\{0\}.$
For $g_{1},g_{2}\in\mathcal{L}^{*}_{k_{1}}(kP)$ and $a\in\mathbb{F}_{q}$, we observe that $g_{1}+g_{2}\in\mathcal{L}(kP)$ and $ag_{1}\in\mathcal{L}^{*}_{k_{1}}(kP)$. It remains to show $g_{1}+g_{2}\notin\mathcal{L}(k_{1}P)$.
Let $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$ be a basis for $\mathcal{L}(kP)$. Since $k_{1}\geq 2g-1$, the space $\mathcal{L}(k_{1}P)$ has a basis contained in $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\}$. Suppose that
$g_{1}+g_{2}=\sum\limits_{i=\ell_{k_{1}}+1}^{\ell_{k}}a_{i}\omega_{i}$
with $a_{i}\in\mathbb{F}_{q}$. It follows that $g_{1}+g_{2}\in\mathcal{L}(k_{1}P)$ implies $g_{1}+g_{2}=a_{1}\omega_{1}+\cdots+a_{\ell_{k_{1}}}\omega_{\ell_{k_{1}}}$. The linear independence of basis elements forces all coefficients to vanish, which means $g_{1}+g_{2}=0\in\mathcal{L}^{*}_{k_{1}}(kP)$. ∎
Theorem IV.1.
Suppose that $F^{\prime}/F$ is a Kummer extension of degree 2 with a totally ramified place $Q_{0}$ in $F^{\prime}/F$, and $Q^{\prime}_{0}|Q_{0}$. Let $s,k,k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0}),f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})\}$
has parameters $[4s,\ell_{k},d_{sr}\geq 2s-k]$.
Proof.
The code $C_{sr,1}$ has length exactly $4s$ over $\mathbb{F}_{q}$. Its dimension is determined by
$k-k_{1}+\ell_{k_{1}}-\dim(\mathcal{L}(kQ_{\infty}-\text{Con}_{F^{\prime}/F}(\sum\limits_{i=1}^{s}Q_{s}))).$
Since $k<2s$, the evaluation map $\text{ev}_{D,1}$ is injective, which implies that the dimension of $C_{sr,1}$ simplifies to $k-k_{1}+\ell_{k_{1}}=\ell_{k}$.
Let $f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0})$ and $f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})$ with $f_{1}\neq 0$. Suppose that $rank(\bar{\varepsilon}_{Q_{i}}(L))=d_{i}$ for $i=1,\ldots,s$. Then the codeword of $L$ has sum-rank distance $d_{L}=\sum\limits_{i=1}^{s}d_{i}$. By Lemma IV.2, we have
$v_{Q_{i}}(\det(\varepsilon_{Q_{i}}(L)))\geq 2-d_{i}$
for each $i$. Furthermore, Lemma IV.1 shows
$\det(\varepsilon_{Q_{i}}(L))\in\mathcal{L}(kQ_{0}).$
From $v_{Q^{\prime}_{0}}(f_{1})<v_{Q^{\prime}_{0}}(f_{2})$ and $f_{1}\neq 0$, we get $\det(\varepsilon_{P_{\alpha}}(L))\neq 0$. When $f_{1}=0$, the matrix $\varepsilon_{Q_{i}}(L)$ takes the form
$A_{\sigma_{2}}\cdot\begin{bmatrix}&f_{21},&f_{22}\\ &f_{22}u,&f_{21}\end{bmatrix}$ (1)
where $f_{2}=f_{21}z_{1}+f_{22}z_{2}$. Then It follows from $\det(\varepsilon_{Q_{i}}(L))=0$ that $u-f_{*}^{2}$ belong to $F$ for a certain function $f_{*}$. Thus $\det(\varepsilon_{Q_{i}}(L))$ cannot vanish as $u\neq\omega^{2}$ for any $\omega\in F$. Therefore $\det(\varepsilon_{P_{\alpha}}(L))$ has at most $k$ zeros. This implies that
$\sum\limits_{i=1}^{s}(2-d_{i})\leq k.$
Consequently, the codeword satisfies $d_{L}\geq 2s-k$. Therefore we have $d_{sr}\geq 2s-k$. This completes the whole proof. ∎
IV-B New Construction for Quadratic Extension
For a quadratic extension $F^{\prime}/F$, we provide an alternative construction by utilizing completely split places. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$. Consider a place $Q_{0}$ that splits completely in $F^{\prime}/F$ with $Q_{0,1}$ and $Q_{0,2}$ lying over $Q_{0}$. Let $G=nQ,G_{i}=nQ_{0,i}$ and $D=Q_{1}+\cdots+Q_{s}$. The evaluation map is constructed as follows:
$\text{ev}_{D,2}:\bigoplus_{i=1}^{2}\mathcal{L}(G_{i})$ $\to\Pi$
$L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$$\mapsto\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))\}$
The image of $\text{ev}_{D,2}$ forms a linear subspace of $\Pi$. Subsequently, the code $C_{sr,2}(D,G)$ is defined as the image. To analyze the parameters of $C_{sr,2}(D,G)$, we require the following lemma.
Lemma IV.4.
Let $Q\neq P$ be a rational divisor of $F$ that splits completely in $F^{\prime}/F$, and $Q_{1},Q_{2}$ the places of $F^{\prime}$ lying over $Q$. For $f_{i}\in\mathcal{L}(kQ_{i})$, we have
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Furthermore, if there exists $f_{i}$ such that
$v_{Q_{i}}(f_{i})<v_{Q_{j}}(f_{j}),\forall j\neq i,$
then $\det(\varepsilon_{P}(L))\neq 0$.
Proof.
First, every term in the determinant expansion of $\varepsilon_{P}(L)$ must be contained in $\mathcal{L}(n(Q_{1}+Q_{2}))$. Note that the conorm of $Q$ is
$\text{Con}_{F^{\prime}/F}(Q)=Q_{1}+Q_{2}.$
Following a similar argument to Lemma IV.1, we obtain
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Second, w.l.o.g, assume that $v_{Q_{1}}(f_{1})<v_{Q_{2}}(f_{2})$. For any non-zero function $g$, we have
$v_{Q_{2}}(\sigma_{2}(g))=v_{Q_{1}}(\sigma_{1}(g)).$
If $L(g)=0$, then
$v_{Q_{1}}(\sigma_{2}(g))=v_{Q_{1}}(f_{1}\sigma_{1}(g))<v_{Q_{1}}(\sigma_{1}(g)).$
This implies
$v_{Q_{2}}(\sigma_{1}(g))<v_{Q_{2}}(\sigma_{2}(g))$
and consequently
$v_{Q_{2}}(f_{1}\sigma_{1}(g))<v_{Q_{2}}(f_{2}\sigma_{2}(g)).$
It follows that $v_{Q_{2}}(L(g))<\infty$, contradicting $L(g)=0$. Hence $L(g)\neq 0$ for all non-zero $g\in\mathcal{O}^{\prime}_{P}$. The conclusion follows from Lemma III.1 (iii). ∎
The subsequent result parallels Theorem IV.1, thus we omit the details of the proof.
Theorem IV.2.
Suppose that $F^{\prime}/F$ is a function field extension of degree 2. Assume that there exists a place $Q_{0}$ of $F$ such that $Q_{0}$ splits completely in $F^{\prime}/F$ with $Q_{0,1}|Q_{0}$ and $Q_{0,2}|Q_{0}$. Let $s,k,k_{1}$ be integers such that $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For $s$ distinct places $Q_{1},\ldots,Q_{s}$, the code
$C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
has parameters $[4s,\ell_{k},d_{sr}\geq 2s-k]$.
IV-C Examples from Elliptic Function Fields
An elliptic function field $\mathcal{E}/\mathbb{F}_{q}$ is a quadratic extension of the rational function field $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ of genus 1. This serves as a canonical example since evaluation places can be selected from all $q+1$ places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$, limiting the code length to $q$. Nevertheless, this suffices to illustrate our framework.
In this section, we consider the case when $q$ is odd, as elliptic function fields over even characteristic are Artin-Schreier extensions. Specifically, let $\mathcal{E}/\mathbb{F}_{q}=\mathbb{F}_{q}(x,y)$ where $x,y$ are transcendental over $\mathbb{F}_{q}$ and satisfy the smooth, irreducible equation:
$y^{2}=f_{\mathcal{E}}$
with $f_{\mathcal{E}}\in\mathbb{F}_{q}[x]$ being a square-free cubic polynomial. The Galois group $\text{Aut}(\mathcal{E}/\mathbb{F}_{q}(x))=\{Id,\sigma\}$ is cyclic of order 2, where
$\sigma(y)=-y.$
Let $P_{\infty}$ be the place at infinity of $\mathbb{F}_{q}(x)$, which is totally ramified in $\mathcal{E}/\mathbb{F}_{q}(x)$. Suppose that $Q_{\infty}$ is the place of $\mathcal{E}$ lying over $P_{\infty}$. The basis $\{1,y\}$ of $\mathcal{E}/\mathbb{F}_{q}(x)$ forms an integral basis for all places of $\mathcal{E}/\mathbb{F}_{q}$ except $P_{\infty}$. Consequently, for any rational place $P\neq P_{\infty}$, the matrix induced by $\sigma$ on $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is
$\begin{bmatrix}1,&0\\ 0,&-1\end{bmatrix}.$
Let $f_{1}$ and $f_{2}$ be two functions over $\mathcal{E}/\mathbb{F}_{q}$, where each $f_{i}$ has the form $f_{i}=f_{i1}+f_{i2}y$ with $f_{i1}$ and $f_{i2}$ belonging to $\mathbb{F}_{q}(x)$ for $i=1,2$. For any rational place $P\neq P_{\infty}$, the matrix induced by $f_{i}$ on $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is
$\begin{bmatrix}f_{i1},&f_{i2}\\ f_{i2}f_{\mathcal{E}},&f_{i1}\end{bmatrix}.$
Let $L=f_{1}+f_{2}\sigma$. Then we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}f_{11}+f_{21},&f_{12}+f_{22}\\ (f_{12}-f_{22})f_{\mathcal{E}},&f_{11}-f_{21}\end{bmatrix}$
and
$\det(\varepsilon_{P}(L))=f_{11}^{2}-f_{21}^{2}-(f_{12}^{2}-f_{22}^{2})f_{\mathcal{E}}.$
For positive integers $k,k_{1}$, define sum-rank codes $C_{sr,1}$ and $C_{sr,2}$ as follows:
$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{\infty}),f_{2}\in\mathcal{L}(k_{1}Q_{\infty})\}$
and
$C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
where $Q_{0,1}$ and $Q_{0,2}$ are rational places lying over $Q_{0}\in\mathbb{F}_{q}(x)/\mathbb{F}_{q}$. Through the Singleton bound combined with Theorems IV.1 and IV.2, we derive
$4s-2k\leq 2d_{sr,1},2d_{sr,2}\leq 4s-k+2,$
where $s\leq q$. The following two examples are given by SageMath [42].
Example IV.1.
Let $q=7$ and consider the elliptic curve $\mathcal{E}:y^{2}=x^{3}+3$ over $\mathbb{F}_{q}$. The rational places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ are
$\{P_{0},P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$
where $P_{\alpha}$ corresponds to the prime element $x-\alpha$. For $k=6$ and $k_{1}=3$, the space $\mathcal{L}(6Q_{\infty})$ has basis $\{1,x,y,x^{2},xy,x^{3}\}$. The code $C_{sr,1}(D,6Q_{\infty})$ has parameters $[24,6,6]$. Taking $f_{1}=3$ and $f_{2}=y$, we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}3,&1\\ -(x^{3}+3),&3\end{bmatrix}$
with
$\det(\varepsilon_{P}(L))=x^{3}+5.$
Since $x^{3}+5$ is irreducible over $\mathbb{F}_{q}[x]$, the function $\det(\varepsilon_{P}(L))$ has no rational zeros. Consequently, the codeword
$\left(\begin{bmatrix}3,&1\\ -3,&3\end{bmatrix},\ldots,\begin{bmatrix}3,&1\\ 5,&3\end{bmatrix}\right)$
achieves the maximum sum-rank distance $14=2s$. Alternatively, taking $f_{1}=1$ and $f_{2}=x^{3}$, we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}x^{3}+1,&0\\ 0,&1-x^{3}\end{bmatrix}$
with
$\det(\varepsilon_{P}(L))=1-x^{6}.$
All places in $\{P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$ are zeros of $\det(\varepsilon_{P}(L))$. Thus, the codeword
$\left(\begin{bmatrix}1,&0\\ 0,&1\end{bmatrix},\ldots,\begin{bmatrix}0,&0\\ 0,&2\end{bmatrix}\right)$
has sum-rank distance 8, attaining the lower bound.
Example IV.2.
Consider the same finite field $\mathbb{F}_{q}$ and elliptic curve $\mathcal{E}$ as in the previous example. We choose the place $P_{1}$, and let $Q_{1,1}$ and $Q_{1,2}$ be the places of $\mathcal{E}/\mathbb{F}_{q}$ lying over $P_{1}$. We take the rational places
$\{P_{0},P_{2},P_{3},P_{4},P_{5},P_{6}\}.$
Let $D^{\prime}=P_{0}+P_{2}+P_{3}+P_{4}+P_{5}P_{6}$. Then the code $C_{sr,2}(D,6P_{1})$ has parameters $[24,6,6]$ with sum-rank weight distribution:
$A_{0}=1,A_{6}=36,A_{7}=144,A_{8}=1542,A_{9}=7944,$
$A_{10}=26904,A_{11}=46959,A_{12}=34122$
where $A_{i}$ counts the number of codewords with sum-rank weight $i$.
V Conclusion
In this paper, we conducted two families of $2\times 2$ sum-rank metric codes by exploiting quadratic Kummer extensions of algebraic function fields. By representing elements in the extension fields as matrices over the base field through explicit algebraic mappings, we rigorously determined the parameters of the proposed codes, including their minimum sum-rank distances and code rates. To demonstrate the practical applicability of our framework, we provided concrete instantiations over elliptic function fields.
To the best of our knowledge, these code constructions are novel in the context of algebraic-geometric (AG) codes for the sum-rank metric. Prior to this work, the only systematic study of AG codes in this metric was established by Berardini and Caruso [4], who proposed a general framework for such codes. However, our methodology differs fundamentally from theirs by explicitly exploiting the intrinsic algebraic structures of function field extensions—particularly through matrix representations derived from quadratic Kummer extensions—to control code parameters and optimize distance properties.
In this paper, we investigate Kummer extensions defined by the equation $y^{2}=u\quad\text{where }u\notin\{\omega^{2}\mid\omega\in F\setminus\mathbb{F}_{q}\}.$ This specific choice of non-square elements $u$ enables us to rigorously establish the non-degeneracy condition $\det(\varepsilon_{P}(L))\neq 0$ in the restricted case where $f_{1}=0$. The principal challenge in generalizing these constructions to $n\times n$ matrices stems from the need to preserve the non-vanishing determinant property $\det(\varepsilon_{P}(L))\neq 0\quad\text{for arbitrary }n\geq 2.$ We intend to address this dimensionality constraint through more flexible algebraic frameworks in subsequent work.
Regarding the sum-rank distance of our AG codes in the sum-rank metric, for a fixed dimension $k-g(F)+1$, we observe the following approximate relationship:
$\max d_{sr}-\min d_{sr}\approx\frac{k+g(F)+1}{2}.$
This raises an interesting problem: whether techniques developed for constructing MDS AG codes [11] can be adapted to the sum-rank metric setting to achieve MSRD AG codes. Following the approach for AG codes with narrow bounds (such as elliptic codes [11]), we propose this problem as an important objective for our subsequent research.
VI Acknowledgment
The work of Chang-An Zhao is supported by Guangdong Basic and Applied Basic Research Foundation(No. 2025A1515011764), Guangdong Major Project of Basic and Applied Basic Research (No. 2019B030302008), the National Natural Science Foundation of China (No. 12441107) and Guangdong Provincial Key Laboratory of Information Security Technology(No. 2023B1212060026).