### v2

# Constructing Sum-Rank Metric Codes from Quadratic Galois Extensions of Function Fields

Abstract
This paper introduces new constructions of sum-rank metric codes derived from algebraic function fields, as existing results on such codes remain limited. A major challenge lies in the determination of their parameters. We address this issue by employing quadratic Galois extensions, proposing two general constructions of $2\times 2$ sum-rank codes. Analogous to algebraic geometry codes in the Hamming metric, our codes achieve a larger block length compared to existing constructions. We determine explicit parameters including dimensions and minimum distances of our codes, and we present an illustrative example using elliptic function fields. Finally, we discuss the asymptotic behavior of our codes and compare them with the Gilbert-Varshamov-like bound for sum-rank metric codes.
Index terms: Sum-rank metric codes, algebraic geometry codes, function fields, elliptic curves.
I Introduction
Algebraic function fields over finite fields provide a fundamental framework for modern error-correcting codes in the Hamming metric [16], as the function field perspective enables a deeper analysis of code parameters, including minimum distance and duality properties [43]. Specifically, a linear $[n,k]$ Hamming metric code $C$ over $\mathbb{F}_{q}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum Hamming distance
$d_{H}:=\min_{c_{1}\neq c_{2}}\{wt_{H}(c_{1}-c_{2})\mid c_{1},c_{2}\in C\},$
where $wt_{H}$ denotes the Hamming weight, counting non-zero coordinate positions. Classical algebraic geometry (AG) codes are constructed by evaluating functions from Riemann-Roch spaces at rational places, enabling explicit determination of code parameters.
For a linear $n,k,d_{H}]$ code $C$, the Singleton bound states that
$d_{H}\leq n-k+1.$
When equality holds, the code is called a maximum distance separable (MDS) code. Reed-Solomon (RS) codes [19], derived from rational function fields $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ with $x$ transcendental over $\mathbb{F}_{q}$, represent prominent examples of Hamming metric MDS codes. Their maximum length of $q+1$ over $\mathbb{F}_{q}$ aligns with the main conjecture of MDS codes.
Recently, sum-rank metric codes have garnered significant attention due to their applications in multishot network coding [24, 26, 33, 37], space-time coding [42], and distributed storage [25, 27]. For further details and applications, we refer the reader to [29]. The basic definitions are provided as follows.
Definition I.1.
Given $s\geq 1$ and positive integers $n_{1},\ldots,n_{s},m_{1},\ldots,m_{s}$ with $n_{i}\leq m_{i}$ and $m_{1}\geq m_{2}\cdots\geq m_{s}$, define
$\Pi:=\Pi_{q}(n_{1}\times m_{1}|\cdots|n_{s}\times m_{s})=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{n_{i}\times m_{i}}$
where $\mathbb{F}_{q}^{n_{i}\times m_{i}}$ denotes the space of $n_{i}\times m_{i}$ matrices over $\mathbb{F}_{q}$. Note that $\Pi$ is an $\mathbb{F}_{q}$-linear space of dimension $\sum_{i=1}^{s}n_{i}m_{i}$. The sum-rank weight of
$X:=(X_{1},\ldots,X_{s})\in\Pi$
is defined as
$wt_{sr}(X):=\sum\limits_{i=1}^{s}{\rm rank}(X_{i}).$
The sum-rank distance between $X$ and $Y$ in $\Pi$ is
$d_{sr}(X,Y)=wt_{sr}(X-Y).$
The sum-rank weight and distance induce a metric on $\Pi$. We present the definition as follows.
Definition I.2.
A linear sum-rank metric code $C_{sr}$ is an $\mathbb{F}_{q}$-linear subspace of $\Pi$. The minimum sum-rank distance of $C_{sr}$ is defined as
$d_{sr}(C):=\min_{X\neq Y}\{wt_{sr}(X-Y)|X,Y\in C_{sr}\}.$
Sum-rank metric codes generalize Hamming metric codes of length $s$ when $n_{i}=m_{i}=1$ for all $1\leq i\leq s$. When $s=1$, they reduce to rank metric codes, which were first introduced by Delsarte [12].
Fundamental properties and bounds for sum-rank metric codes are examined in [1, 2, 5, 39]. The Singleton bound was established in [5] and [27]. In the general case, let $j$ and $\delta$ be the unique integers satisfying $d-1=\sum_{i=1}^{j-1}n_{i}+\delta$ with $0\leq\delta\leq n_{j}-1$. Then
$|C|\leq q^{\sum_{i=j}^{s}m_{i}n_{i}-m_{j}\delta}.$
A code $C\subseteq\Pi$ that attains this bound is called a maximum sum-rank distance (MSRD) code. For the special case where $n_{1}=\cdots=n_{s}=n$ and $m_{1}=\cdots=m_{s}=m$, this bound takes the form
$|C|\leq q^{m(sn-d_{sr}+1)}.$
Throughout this paper, we concentrate on sum-rank metric codes with these uniform parameters.
Let $N:=sn$. For fixed positive real numbers $\mathcal{R}_{sr},\delta_{sr}\in(0,1)$ satisfying
$\mathcal{R}_{sr}<$ $\delta_{sr}^{2}-\delta_{sr}(2+\frac{2}{N})+1+\frac{2}{N}+\frac{1}{N^{2}}$
$-\frac{\sum_{i=1}^{d_{sr}N-1}\log_{q}(1+\frac{s-1}{i})+\log_{q}(\delta_{sr}N-1)}{Nm}-\frac{\log_{q}(\gamma_{q})}{nm},$
where $\gamma_{q}=\prod_{i=1}^{\infty}(1-q^{-i})^{-1}$, there exists a sum-rank code with rate at least $\mathcal{R}_{sr}$ and relative minimum sum-rank distance at least $\delta_{sr}$. When $m=\xi n$ goes to the infinity and $m\in\omega(\log_{q}(s))$, where $\xi$ is a constant, we have
$\mathcal{R}_{sr}\sim\delta_{sr}^{2}-\delta_{sr}(1+\frac{1}{\xi})+1.$
Recent work [39] demonstrates that random linear sum-rank metric codes attain the asymptotic Gilbert-Varshamov-like (GV-like) bound with high probability. Consider the case where $n=m$ and the number of blocks $s$ goes to infinity. Then the previous bound is asymptotically provided by [4]:
$\mathcal{R}_{sr}<$ $(\delta_{sr}-1)^{2}-\frac{\delta_{sr}}{n}\log_{q}(1+\frac{1}{\delta_{sr}n})-\frac{\log_{q}(1+\delta_{sr}n)}{n^{2}}$
$-\frac{\log_{q}(\gamma_{q})}{n^{2}}+o(1).$
Constructions of “good” sum-rank metric codes are available in several works [1, 6, 7, 9, 25, 28, 30, 31, 32, 35, 36]. A standard technique involves constructing $C$ as a classical Hamming metric code of length $sn$ over $\mathbb{F}_{q^{m}}$ and then applies rank metric code constructions to each $n$-length block. This ensures that the dimension $\kappa$ of $C$ over $\mathbb{F}_{q^{m}}$ satisfies $d_{sr}+\kappa\leq sn+1$, matching the Singleton bound for the Hamming metric. Moreover, algorithms over $\mathbb{F}_{q^{m}}$ can be employed for decoding sum-rank metric codes, see [3, 10, 17, 41].
One of the most prominent rank metric codes is the Gabidulin code, which is constructed using linearized polynomials over $\mathbb{F}_{q^{m}}$, see [13, 15, 21, 23, 34]. Among sum-rank metric codes, linearized Reed-Solomon codes [25] represent the most extensively investigated family to date, see [8, 18, 22, 35, 40]. Gabidulin codes and linearized RS codes are maximum rank distance (MRD) and MSRD respectively, serving as generalizations of RS codes.
Motivations
While non-trivial algebraic geometric constructions for sum-rank metric codes exist, see [9, 30], most results are constructed using towers of function fields as illustrated below. In [4], E. Berardini and X. Caruso established the first geometric construction of sum-rank metric codes. Their approach employs the Ore polynomial ring [38] over an extension field derived from the completion of $F$ at specified places, and subsequently develops the theory of Riemann-Roch spaces over this ring. This framework allows them to define linearized algebraic geometry codes and present explicit constructions of sum-rank metric codes with parameters consistent with classical algebraic geometry codes.
A principal motivation for constructing codes via algebraic function fields lies in the availability of numerous evaluation places bounded by the Hasse-Weil bound for higher genus. Consequently, codes with long lengths can be constructed over small finite fields. Classical algebraic geometry codes are known to asymptotically surpass the Gilbert-Varshamov (GV) bound [45]. This naturally raises the question: whether sum-rank codes derived from algebraic geometry also possess parameters surpassing the GV bound.
[…]
Figure 1: Tower of function field extensions.
Our contributions
In contrast to the construction in [4], we propose an approach for constructing sum-rank metric codes using function fields. The central element of our construction is the local integral basis at a place in function field extensions. Specifically, let $\mathcal{O}_{P}$ be the valuation ring at a place $P$ of $F/\mathbb{F}_{q}$, and let $F^{\prime}/F$ be a finite separable function field extension. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ constitutes a vector space over $\mathcal{O}_{P}$. Consequently, a function in $F^{\prime}/\mathbb{F}^{\prime}_{q}$ that is regular at the places lying over $P$ corresponds to a matrix over $F$, where $\mathbb{F}^{\prime}_{q}$ is a finite extension of $\mathbb{F}_{q}$. When $F^{\prime}/F$ is Galois, we define a map on $\mathcal{O}^{\prime}_{P}$, via the direct sum of Riemann-Roch spaces. This yields sum-rank metric codes through the evaluation of the function matrix at rational places. Although our construction shares some similarities with [4], it significantly differs by focusing on the intrinsic structure of function field extensions. By decomposing the Riemann-Roch spaces as a direct sum of subspaces, we explicitly determine the parameters of our sum-rank codes. Furthermore, when $q$ is a special square, we analyze the asymptotic behavior of our codes and compare them with the GV-like bound.
Organization
The remainder of this paper is organized as follows. Section II introduces the fundamental concepts required for our purpose. In Section III, we develop a general construction of sum-rank metric codes via function field extensions. Section IV provides the explicit sum-rank metric codes using quadratic Galois extensions and presents a new specialized construction, including an instance involving elliptic function fields and the comparison with the GV-like bound. Section V concludes this paper.
II Preliminaries and basic concepts
This section reviews fundamental definitions and notations, primarily based on [43].
II-A Places and divisors
Let $q$ be a prime power. An algebraic function field $F/\mathbb{F}_{q}$ is a finite extension of $\mathbb{F}_{q}(x)$ with genus $g(F)$, where $x$ is transcendental over $\mathbb{F}_{q}$. A place $P$ is the maximal ideal of a valuation ring $\mathcal{O}_{P}$ in $F/\mathbb{F}_{q}$. Denote by $\hat{F}_{P}:=\mathcal{O}_{P}/P$ the residue field of $P$, which is a finite extension of $\mathbb{F}_{q}$. The degree of $P$ is defined as
$\deg(P)=[\hat{F}_{P}:\mathbb{F}_{q}].$
If $\deg(P)=1$, then $P$ is called rational. The prime element (or uniformizing parameter) of $P$ is a function $t\in P$ such that $P=t\mathcal{O}_{P}$. Any non-zero element $z\in F$ admits a unique representation $z=t^{n}u$ with $u\in\mathcal{O}^{*}_{P}$. The discrete valuation associated with $P$ is
$v_{P}:F\to\mathbb{Z}\cup\{\infty\}$
, satisfying $v_{P}(z)=n$ and $v_{P}(0)=\infty$. The strict triangle inequality is characterized as follows:
Lemma II.1.
For any discrete valuation $v_{P}$ with associated place $P$,
$v_{P}(x+y)\geq\min\{v_{P}(x),v_{P}(y)\}$
holds for all $x,y\in F$. Moreover, if $v_{P}(x)\neq v_{P}(y)$, then $v_{P}(x+y)=\min\{v_{P}(x),v_{P}(y)\}$.
A divisor $G$ is defined as a formal sum of places
$G:=\sum\limits_{P}n_{P}P$
with $n_{P}\in\mathbb{Z}$ and $\deg(G)=\sum\limits_{P}n_{P}\deg(P)$. The support of $G$ is defined as
$\text{supp}(G):=\{P\mid n_{P}\neq 0\}.$
For any function $z\in F$,
$(z):=\sum\limits_{P}v_{P}(z)P$
is called the principal divisor of $z$. Given two divisors $G_{1}$ and $G_{2}$, we write $G_{1}\leq G_{2}$ if $n_{P}(G_{1})\leq n_{P}(G_{2})$ for all places $P$. The Riemann-Roch space associated with $G$ is defined as
$\mathcal{L}(G):=\{z\in F\mid(z)+G\geq 0\}\cup\{0\}.$
This space is a vector space over $\mathbb{F}_{q}$ with dimension given by the Riemann-Roch theorem:
Lemma II.2.
[43, Theorem 1.4.17] Let $\ell(G)$ denote the dimension of $\mathcal{L}(G)$ over $\mathbb{F}_{q}$. Then
$\ell(G)\geq\deg(G)-g(F)+1.$
Moreover, if $\deg(G)\geq 2g(F)-1$, then
$\ell(G)=\deg(G)-g(F)+1.$
II-B Extensions of function fields
An algebraic function field $F^{\prime}/\mathbb{F}_{q^{m}}$ is called an algebraic extension of $F/\mathbb{F}_{q}$ if $F^{\prime}/F$ is an algebraic field extension with $F^{\prime}:F]<\infty$ and $m\geq 1$. A place $P^{\prime}$ of $F^{\prime}$ is said lying over a place $P$ of $F$ if $P\subseteq P^{\prime}$, denoted by $P^{\prime}|P$. The ramification index $e(P^{\prime}|P)$ is defined as the integer satisfying
$v_{P^{\prime}}(x)=e(P^{\prime}|P)\cdot v_{P}(x)$
for all $x\in F$, and the relative degree of $P^{\prime}$ over $P$ is given by
$f(P^{\prime}|P)=[\mathcal{O}_{P^{\prime}}/P^{\prime}:\mathcal{O}_{P}/P].$
For any place $P$ of $F$, there exist only finitely many places $P^{\prime}$ of $F^{\prime}$ such that $P^{\prime}|P$. The conorm of $P$ is defined as
$\text{Con}_{F^{\prime}/F}(P):=\sum\limits_{P^{\prime}|P}e(P^{\prime}|P)\cdot P^{\prime}$
and
$\text{Con}_{F^{\prime}/F}\left(\sum\limits_{P}n_{P}P\right):=\sum n_{P}\cdot\text{Con}_{F^{\prime}/F}(P).$
Consequently, for any non-zero function $z\in F$, the principal divisor of $z$ in $F^{\prime}$ satisfies
$(z)_{F^{\prime}}=\text{Con}_{F^{\prime}/F}((z)_{F}).$
The fundamental equality of places is stated as follows.
Proposition II.1.
[43, Theorem 3.3.11] Let $P$ be a place of $F$, and let $P_{1},\ldots,P_{n}$ be all places of $F^{\prime}$ lying over $P$. Denote $e_{i}=e(P_{i}|P)$ and $f_{i}=f(P_{i}|P)$ for $i=1,\ldots,n$. Then
$\sum\limits_{i=1}^{n}e_{i}f_{i}=[F^{\prime}:F].$
If $F^{\prime}:F]=n$ and there exist exactly $n$ distinct places lying over $P$, then $P$ splits completely in $F^{\prime}/F$. In this case, $e_{i}=f_{i}=1$ for all $i$. If $e(P^{\prime}|P)>1$ for some $P^{\prime}$, then $P$ is ramified in $F^{\prime}/F$. Moreover, if there exists a place $P^{\prime}$ with $e(P^{\prime}|P)=n$, then $P$ is totally ramified in $F^{\prime}/F$.
II-C Galois Extensions
An extension $F^{\prime}/\mathbb{F}_{q^{m}}$ of a function field $F/\mathbb{F}_{q}$ is Galois if $F^{\prime}/F$ is a finite Galois extension. Specifically, the automorphism group
[…]
has order $F^{\prime}:F]$. This group is called the Galois group and write $\text{Gal}(F^{\prime}/F):=\text{Aut}(F^{\prime}/F)$.
For any place $P$ of $F/\mathbb{F}_{q}$, the Galois group $\text{Gal}(F^{\prime}/F)$ acts transitively on the set of places
[…]
via $\sigma(P^{\prime})=\{\sigma(x)\mid x\in P^{\prime}\}$. For any place $P^{\prime}$ of $F^{\prime}/\mathbb{F}_{q^{m}}$, the valuation corresponding to the place $\sigma(P^{\prime})$ is given by
$v_{\sigma(P^{\prime})}(y)=v_{P^{\prime}}(\sigma^{-1}(y)),\quad y\in F^{\prime}.$
III A General Setting of Sum-Rank Metric Codes
In this section, we introduce a general construction of sum-rank metric codes using function field extensions, which differs from the approach proposed by E. Berardini and X. Caruso in [4].
III-A The Main Idea of Matrices over $\mathbb{F}_{q}$
Let $F/\mathbb{F}_{q}$ be a function field, and $F^{\prime}/\mathbb{F}_{q^{m}}$ be a Galois extension of $F/\mathbb{F}_{q}$ with $F^{\prime}:F]=n$. Suppose that $P$ is a place of $F$, and $t\in F$ is a prime element of $P$. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ is given by
$\mathcal{O}^{\prime}_{P}=\bigcap\limits_{i=1}^{n}\mathcal{O}_{P_{i}}.$
An integral basis $\{z_{1},\ldots,z_{n}\}$ of $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is a basis of $F^{\prime}/F$ satisfying
$\mathcal{O}^{\prime}_{P}=\sum\limits_{i=1}^{n}\mathcal{O}_{P}z_{i}.$
The quotient $V:=\mathcal{O}^{\prime}_{P}/t\mathcal{O}^{\prime}_{P}$ forms a vector space over the finite field $\hat{F}_{P}$ with basis
$\{z_{1}+t\mathcal{O}^{\prime}_{P},\ldots,z_{n}+t\mathcal{O}^{\prime}_{P}\}.$
For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and a place $P^{\prime}$ with $P^{\prime}|P$, observe that for any $z\in\mathcal{O}^{\prime}_{P}$,
$v_{P^{\prime}}(\sigma(z))=v_{\sigma^{-1}(P^{\prime})}(z)\geq 0$
since $\sigma^{-1}(P^{\prime})|P$. This implies $\sigma(z)\in\mathcal{O}^{\prime}_{P}$, and hence $\sigma(V)=V$. The set
$\{\sigma(z_{1})+t\mathcal{O}^{\prime}_{P},\ldots,\sigma(z_{n})+t\mathcal{O}^{\prime}_{P}\}$
forms another basis of $V$ over $\hat{F}_{P}$. Then $\sigma$ induces an $n\times n$ matrix $\bar{A}_{\sigma}$ over $\hat{F}_{P}$ with respect to the basis $\{z_{i}\}$.
Similarly, for any $f\in\mathcal{O}^{\prime}_{P}$, consider the multiplication map $\mu_{f}:\mathcal{O}^{\prime}_{P}\to\mathcal{O}^{\prime}_{P}$ defined by $\mu_{f}(z)=fz$. This induces an $\hat{F}_{P}$-linear map $\bar{\mu}_{f}:V\to V$ via
$\bar{\mu}_{f}(z+t\mathcal{O}^{\prime}_{P}):=fz+t\mathcal{O}^{\prime}_{P}.$
This also induces an $n\times n$ matrix $\bar{B}_{f}$ with entries in $\hat{F}_{P}$.
Remark III.1.
The matrix $A=(a_{ij})$ associated with $\sigma$ has entries in $\mathcal{O}_{P}$ due to the integral basis property. Under the reduction map
$\pi:\mathcal{O}_{P}\to\hat{F}_{P},$
we obtain $\bar{A}=(\pi(a_{ij}))$. This justifies the notation $\bar{A}_{\sigma}$ and $\bar{B}_{f}$.
When $P$ is rational (i.e., $\hat{F}_{P}=\mathbb{F}_{q}$), both $\sigma\in\text{Gal}(F^{\prime}/F)$ and $f\in\mathcal{O}^{\prime}_{P}$ induce matrices over $\mathbb{F}_{q}$. Let $\text{Gal}(F^{\prime}/F)=\{\sigma_{1},\ldots,\sigma_{n}\}$, and consider functions $f_{1},\ldots,f_{n}\in\mathcal{O}^{\prime}_{P}$. We define a linear operator
$L:\mathcal{O}^{\prime}_{P}$ $\to\mathcal{O}^{\prime}_{P}$
$g$ $\mapsto\sum\limits_{i=1}^{n}f_{i}\sigma_{i}(g).$
For any two functions $g_{1},g_{2}\in F$, we have
$L(g_{1}+g_{2})=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}(g_{1}+g_{2})=\sum\limits_{i=1}^{n}f_{i}(g_{1}+g_{2})=L(g_{1})+L(g_{2}).$
Thus, this $F^{\prime}$-linear operator naturally induces a linear morphism on $F$. We further define a map $\varepsilon_{P}$ such that
$\varepsilon_{P}(L)=\sum\limits_{i=1}^{n}A_{i}B_{i}=(M_{ij})$
where $B_{i}$ and $A_{i}$ are matrices representing $\mu_{f_{i}}$ and $\sigma_{i}$ for $1\leq i\leq n$, respectively. Then $\varepsilon_{P}(L)$ is an $n\times n$ matrix over $\mathcal{O}_{P}$. Moreover, let
$\bar{\varepsilon}_{P}(L)=\sum\limits_{i=1}^{n}\bar{A_{i}}\bar{B_{i}}=(\pi(M_{ij}))$
be the matrix over $\mathbb{F}_{q}$ induced by $\varepsilon_{P}(L)$. We now state the following lemma.
Lemma III.1.
  • (i)
    Let $\{z_{1},\ldots,z_{n}\}$ be an integral basis of $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$. Then the matrix representing $L$ with respect to this basis is $\varepsilon_{P}(L)$.
  • (ii)
    The determinant $\det(\varepsilon_{P}(L))$ belongs to $\mathcal{O}_{P}$.
  • (iii)
    The kernel ${\rm ker}(L)\neq\{0\}$ if and only if $\det(\varepsilon_{P}(L))=0$.
Proof.
(i) Let $g=g_{1}z_{1}+\cdots+g_{n}z_{n}\in\mathcal{O}^{\prime}_{P}$. For any $\sigma\in\text{Gal}(F^{\prime}/F)$ and $f\in\mathcal{O}_{P}$, let $A_{\sigma}$ and $B_{f}$ denote the matrices defined above. Then
$\mu_{f}(g)=[g_{1},\ldots,g_{n}]B_{f}[z_{1},\ldots,z_{n}]^{T}$
and
$\sigma(g)=[g_{1},\ldots,g_{n}]A_{\sigma}[z_{1},\ldots,z_{n}]^{T}.$
Thus $f\sigma(g)=[g_{1},\ldots,g_{n}]B_{f}A_{\sigma}[z_{1},\ldots,z_{n}]^{T}$. The result follows from the linearity of the maps $\mu_{f}$ and $\sigma$.
(ii) The result immediately follows since all coefficients of $A_{\sigma_{i}}$ and $B_{f_{i}}$ are contained in $\mathcal{O}_{P}$.
(iii) Consider the equation
$\varepsilon_{P}(L)^{T}[x_{1},\ldots,x_{n}]^{T}=[0,\ldots,0].$
Any solution $g_{1},\ldots,g_{n}]$ satisfies $L(g_{1}z_{1}+\cdots+g_{n}z_{n})=0$. Therefore ${\rm ker}(L)$ contains nonzero elements if and only if the system has nontrivial solutions, which occurs precisely when $\det(\varepsilon_{P}(L))=0$. ∎
III-B Sum-rank Codes in the General Case
We now present the construction of sum-rank metric codes. Let $Q_{1},\ldots,Q_{s}$ be distinct rational places of $F$, and $Q_{0}$ be another place of $F$. Assume that $P_{0,1},\ldots,P_{0,m}$ are all places of $F^{\prime}$ lying over $Q_{0}$.Let $G=nQ_{0}$, $D=Q_{1}+\cdots+Q_{s}$, and $G_{j}$ are $n$ divisor of $F^{\prime}$ with
$\text{supp}(G_{j})\subset\{P_{0,1},\ldots,P_{0,m}\}.$
Define the evaluation map:
$\text{ev}_{D}:\mathcal{L}(G_{1})\oplus\cdots\oplus\mathcal{L}(G_{n})$ $\to\Pi=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{n\times n}$
$(f_{1},\ldots,f_{n})$ $\mapsto(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))$
where $L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$. Since $f_{1},\ldots,f_{n}$ have poles only at $P_{0}$, they belong to $\mathcal{O}^{\prime}_{Q_{i}}$ for all $1\leq i\leq s$. Thus $\varepsilon_{Q_{i}}(L)$ are well-defined. Note that $\mathcal{L}(G_{j})$ is an $\mathbb{F}_{q}$-linear space. Since $L$ is $F$-linear, it follows that $\text{ev}_{D}$ is an $\mathbb{F}_{q}$-linear map. Thus the image of $\text{ev}_{D}$ is a subspace of $\Pi$. The code $C_{sr}(D,G)$ is defined as the image of $\text{ev}_{D}$, which is exactly a linear space over $\mathbb{F}_{q}$.
Remark III.2.
In the work of [4], E. Berardini and X. Caruso constructed codes through the isomorphism chain
$D_{F^{\prime}_{i},x}\to D_{F^{\prime}_{i},1}\to\text{End}_{F_{i}}(F^{\prime}_{i})$
where $F_{i}$ denotes the completion of $F$ at $Q_{i}$, and $F^{\prime}_{i}$ denotes the product of completions of $F^{\prime}$ at places lying over $Q_{i}$ for each $i=1,\ldots,s$. The functor for generating the matrix in [4] has form $\sum\limits_{i=0}^{n-1}f_{i}\Phi^{i}$, where $\Phi^{r}=x$ for some $0\neq x\in F$, and $f_{i}\in\Lambda_{L,x}(G^{\prime})$. To determine the code parameters, E. Berardini and X. Caruso hypothesized that $D_{L,x}$ has no nonzero zero divisor.
Our approach directly employs the operator $L=\sum\limits_{i=1}^{n}f_{i}\sigma_{i}$, corresponding to elements of $D_{F^{\prime},1}$ since $F^{\prime}/F$ is Galois. However, the ring $D_{F^{\prime},1}$ is not a domain because the polynomial $T^{r}-1\in F^{\prime}[T]$ is reducible. For the determination of code parameters, we consider $Q_{0}$ that is totally ramified or splits completely in $F^{\prime}/F$, and then present another choice for $f_{i}$, which is explicitly described in subsequent sections.
Proposition III.1.
The length $n^{2}s$ of $C_{sr,1}$ over $\mathbb{F}_{q}$ satisfies
$n^{2}s\leq n^{2}(q+2g(F)\sqrt{q}+1).$
Proof.
In the construction of $C_{sr}$, each place $Q\neq Q_{0}$ yields a well-defined $n\times n$ matrix $\bar{\varepsilon}_{P}(L)$ over $\mathbb{F}_{q}$. By the Hasse-Weil bound, the function field $F/\mathbb{F}_{q}$ has at most $q+1+2g(F)\sqrt{q}$ rational places, which implies that $s\leq q+2g(F)\sqrt{q}+1$. ∎
IV Sum-Rank Metric Codes via Quadratic Extension of Algebraic Function Fields
IV-A Construction with Totally Ramified Rational Places
We first establish some necessary lemmas before deriving the dimension and minimum sum-rank distance.
Lemma IV.1.
Let $P$ be a rational place of $F$. Suppose that $Q\neq P$ is a place of $F$ such that $Q$ is totally ramified in $F^{\prime}/F$. Let $Q^{\prime}$ be the place of $F^{\prime}$ lying over $Q$. For $f_{1},\ldots,f_{n}\in\mathcal{L}(kQ^{\prime})$, we have
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Furthermore, if there exists an $f_{i}$ satisfying
$v_{Q^{\prime}}(f_{i})<v_{Q^{\prime}}(f_{j}),\forall j\neq i,$
then $\det(\varepsilon_{P}(L))\neq 0$.
Proof.
Since $f_{i}\in\mathcal{L}(kQ^{\prime})$, the coefficients of $B_{f_{i}}$ belong to $\mathcal{L}(kQ^{\prime})$ for $i=1,\ldots,n$. Therefore, the discrete valuation of $\det(\varepsilon_{P}(L))$ at $Q^{\prime}$ is greater than the minimal valuation in its determinant expansion. Consequently, we have $\det(\varepsilon_{P}(L))\in\mathcal{L}(knQ^{\prime})$. As $Q$ is totally ramified, the conorm of $Q$ is
$\text{Con}_{F^{\prime}/F}(Q)=nQ^{\prime}.$
By Lemma III.1 (ii) and the properties of the conorm, we obtain
$v_{Q}(\det(\varepsilon_{P}(L)))\leq k.$
Note that for any place $Q^{\prime}$ of $F^{\prime}$, there is a unique place $Q$ in $F$ such that $Q^{\prime}|Q$, so $\det(\varepsilon_{P}(L))$ has poles only at $Q$. Hence, $\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ)$. For any nonzero $g\in\mathcal{O}^{\prime}_{P}$, the equality
$v_{Q^{\prime}}(\sigma_{i}(g))=v_{Q^{\prime}}(\sigma_{1}(g))$
holds for all $i=1,\ldots,n$. W.l.o.g, assume that $v_{Q^{\prime}}(f_{1})<v_{Q^{\prime}}(f_{j})$ for all $j=2,\ldots,n$. This implies that
$v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(f_{j}\sigma_{j}(g)).$
By the strict triangle inequality, it follows that
$v_{Q^{\prime}}\left(\sum\limits_{j=1}^{n}f_{j}\sigma_{j}(g)\right)=v_{Q^{\prime}}(f_{1}\sigma_{1}(g))<v_{Q^{\prime}}(\sigma_{1}(g)).$
Thus $L(g)\neq 0$ for any $g\neq 0$. The conclusion follows from Lemma III.1 (iii). ∎
Lemma IV.2.
Let $P$ be a place of $F$ with a prime element $t$. If $\det(\varepsilon_{P}(L))\neq 0$ for $L$ associated to $Q\neq P$ while ${\rm rank}(\bar{\varepsilon}_{P}(L))<n$, then $\det(\varepsilon_{P}(L))\in P$.
Proof.
The condition ${\rm rank}(\bar{\varepsilon}_{P}(L))<n$ means that $\det(\bar{\varepsilon}_{P}(L))=0$. Let $\pi_{P}$ denote the residue class map at $P$. For $\varepsilon_{P}(L)=(M_{ij})$, we have
$\bar{\varepsilon}_{P}(L)=(\pi_{P}(M_{ij}))=(M_{ij})(P).$
Therefore
$\det(\bar{\varepsilon}_{P}(L))=\det(\varepsilon_{P}(L))(P).$
Since $\det(\varepsilon_{P}(L))(P)=0$, it follows that $\det(\varepsilon_{P}(L))\in P$. ∎
Let $k,k_{1}$ be integers such that $k>k_{1}\geq 2g(F)-1$, and $P$ be a rational place of $F$. Then by the Riemann-Roch theorem, we obtain $\mathcal{L}(kP)$ and $\mathcal{L}(k_{1}P)$ are vector spaces over $\mathbb{F}_{q}$ with dimensions
$\ell_{k}:=k-g(F)+1,\ell_{k_{1}}:=k_{1}-g(F)+1,$
respectively. Let
$\mathcal{L}(kP)=\langle\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k}}\rangle$
where $\{\omega_{1},\omega_{2},\ldots,\omega_{\ell_{k_{1}}}\}$ forms a basis for $\mathcal{L}(k_{1}P)$. We define
$\mathcal{L}^{*}_{k_{1}}(kP):=\langle\omega_{\ell_{k_{1}}+1},\ldots,\omega_{\ell_{k}}\rangle.$
It is obvious that $\mathcal{L}^{*}_{k_{1}}(kP)$ is a vector space over $\mathbb{F}_{q}$ of dimension $k-k_{1}$. Then we have:
Theorem IV.1.
Suppose that $F^{\prime}/F$ is a Galois extension of degree 2 with a totally ramified rational place $Q_{0}$ in $F^{\prime}/F$, and $Q^{\prime}_{0}|Q_{0}$. Let $s,k$ and $k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For distinct places $Q_{1},\ldots,Q_{s}$, the code
$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0}),f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})\}$
has parameters $4s,\ell_{k},d_{sr}\geq 2s-k]$.
Proof.
The length of $C_{sr,1}$ is exactly $4s$ over $\mathbb{F}_{q}$. Its dimension is determined by
$k-k_{1}+\ell_{k_{1}}-\dim(\mathcal{L}(kQ_{\infty}-\text{Con}_{F^{\prime}/F}(\sum\limits_{i=1}^{s}Q_{s}))).$
Since $k<2s$, the evaluation map $\text{ev}_{D,1}$ is injective, implying that the dimension of $C_{sr,1}$ reduces to $k-k_{1}+\ell_{k_{1}}=\ell_{k}$.
Let $f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ^{\prime}_{0})$ and $f_{2}\in\mathcal{L}(k_{1}Q^{\prime}_{0})$ with $f_{1}\neq 0$. Suppose that $rank(\bar{\varepsilon}_{Q_{i}}(L))=d_{i}$ for $i=1,\ldots,s$. Then the sum-rank weight of the codeword corresponding to $L$ is $d_{L}=\sum\limits_{i=1}^{s}d_{i}$. By Lemma IV.2, we have
$v_{Q_{i}}(\det(\varepsilon_{Q_{i}}(L)))\geq 2-d_{i}$
for each $i$. Furthermore, Lemma IV.1 shows that
$\det(\varepsilon_{Q_{i}}(L))\in\mathcal{L}(kQ_{0}).$
From $v_{Q^{\prime}_{0}}(f_{1})<v_{Q^{\prime}_{0}}(f_{2})$ and $f_{1}\neq 0$, it follows that $\det(\varepsilon_{Q_{i}}(L))\neq 0$.
Denote $F^{\prime}=F(y)$. Since $F^{\prime}:F]=2$, there exists a monic irreducible quadratic polynomial $\phi(T)\in F[T]$ that is the minimal polynomial of $y$ over $F$, write $\phi(T)=T^{2}+a_{1}T+a_{2}$.
When $f_{1}=0$, the matrix $\varepsilon_{Q_{i}}(L)$ takes the form
$A_{\sigma_{2}}\cdot\begin{bmatrix}&f_{21},&f_{22}\\ &-f_{22}a_{2},&f_{21}-f_{22}a_{1}\end{bmatrix}$
where $f_{2}=f_{21}z_{1}+f_{22}z_{2}$. Then $\det(\varepsilon_{Q_{i}}(L))=0$ means
$f_{21}^{2}-f_{21}f_{22}a_{1}+f_{22}^{2}a_{2}=0.$
If $f_{22}=0$, the previous equation implies $f_{21}=0$ and $f_{2}=0$. If $f_{21}\neq 0$, then the function $\frac{-f_{21}}{f_{22}}$ is a root of $\phi(T)$. Both cases lead to a contradiction. Thus $\det(\varepsilon_{Q_{i}}(L))$ cannot vanish, and it has at most $k$ zeros. This implies that
$\sum\limits_{i=1}^{s}(2-d_{i})\leq k.$
Consequently, the codeword satisfies $d_{L}\geq 2s-k$. Therefore, we have $d_{sr}\geq 2s-k$. ∎
IV-B Construction with Splitting Completely Rational Places
For a Galois extension $F^{\prime}/F$ of degree 2, the following lemmas are required to determine code parameters.
Lemma IV.3.
Let $P$ be a rational place of $F$. Suppose that $Q\neq P$ is a place of $F$ that splits completely in $F^{\prime}/F$, and $Q_{1},Q_{2}$ are the places of $F^{\prime}$ lying over $Q$. For $f_{i}\in\mathcal{L}(kQ_{i})$, we have
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Furthermore, if there exists an $f_{i}$ satisfying
$v_{Q_{i}}(f_{i})<v_{Q_{j}}(f_{j}),\forall j\neq i,$
then $\det(\varepsilon_{P}(L))\neq 0$.
Proof.
First, every term in the determinant expansion of $\varepsilon_{P}(L)$ is contained in $\mathcal{L}(n(Q_{1}+Q_{2}))$. Note that the conorm of $Q$ is
$\text{Con}_{F^{\prime}/F}(Q)=Q_{1}+Q_{2}.$
By an argument similar to that in Lemma IV.1, we obtain
$\det(\varepsilon_{P}(L))\in\mathcal{L}(kQ).$
Second, w.l.o.g, assume that $v_{Q_{1}}(f_{1})<v_{Q_{2}}(f_{2})$. For any non-zero function $g$, we have
$v_{Q_{2}}(\sigma_{2}(g))=v_{Q_{1}}(\sigma_{1}(g)).$
If $L(g)=0$, that is, $f_{1}\sigma_{1}(g)+f_{2}\sigma_{2}(g)=0$. Then by the strict triangle inequality and $v_{Q_{1}}(f_{1})\leq 0$, it follows that
$v_{Q_{1}}(\sigma_{2}(g))=v_{Q_{1}}(f_{1}\sigma_{1}(g))\leq v_{Q_{1}}(\sigma_{1}(g)).$
This implies
$v_{Q_{2}}(\sigma_{1}(g))\leq v_{Q_{2}}(\sigma_{2}(g))$
and consequently
$v_{Q_{2}}(f_{1}\sigma_{1}(g))<v_{Q_{2}}(f_{2}\sigma_{2}(g)).$
It follows that $v_{Q_{2}}(L(g))<\infty$, contradicting $L(g)=0$. Hence $L(g)\neq 0$ for all non-zero $g\in\mathcal{O}^{\prime}_{P}$. The conclusion follows from Lemma III.1 (iii). ∎
The subsequent result is analogous to Theorem IV.1, so we omit proof details.
Theorem IV.2.
Suppose that $F^{\prime}/F$ is a function field extension of degree 2. Assume that there exists a place $Q_{0}$ of $F$ such that $Q_{0}$ splits completely in $F^{\prime}/F$ with $Q_{0,1},Q_{0,2}$ lying over $Q_{0}$. Let $s,k$ and $k_{1}$ be integers satisfying $s\geq 2$ and $2g(F)-1\leq k_{1}<k<2s$. For distinct places $Q_{1},\ldots,Q_{s}$, the code
$C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}\sigma_{1}+f_{2}\sigma_{2},f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
has parameters $4s,\ell_{k},d_{sr}\geq 2s-k]$.
IV-C Examples from Elliptic Function Fields
An elliptic function field $\mathcal{E}/\mathbb{F}_{q}$ is a quadratic extension of the rational function field $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ and has genus 1. It serves as a canonical example and suffices to illustrate our framework.
In this section, we consider the case when $q$ is odd. Specifically, let $\mathcal{E}/\mathbb{F}_{q}=\mathbb{F}_{q}(x,y)$ where $x,y$ are transcendental over $\mathbb{F}_{q}$ and satisfy the smooth, irreducible equation:
$y^{2}=f_{\mathcal{E}}$
where $f_{\mathcal{E}}\in\mathbb{F}_{q}[x]$ is a square-free cubic polynomial. The Galois group $\text{Aut}(\mathcal{E}/\mathbb{F}_{q}(x))=\{Id,\sigma\}$ is cyclic of order 2, where
$\sigma(y)=-y.$
Let $P_{\infty}$ be the place at infinity of $\mathbb{F}_{q}(x)$, which is totally ramified in $\mathcal{E}/\mathbb{F}_{q}(x)$. Suppose that $Q_{\infty}$ is the place of $\mathcal{E}$ lying over $P_{\infty}$. The basis $\{1,y\}$ of $\mathcal{E}/\mathbb{F}_{q}(x)$ forms an integral basis for all places of $\mathcal{E}/\mathbb{F}_{q}$ except $P_{\infty}$. Consequently, for any rational place $P\neq P_{\infty}$, the matrix induced by $\sigma$ on $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is
$\begin{bmatrix}1,&0\\ 0,&-1\end{bmatrix}$.
Let $f_{1}$ and $f_{2}$ be two functions over $\mathcal{E}/\mathbb{F}_{q}$, where each $f_{i}$ has the form $f_{i}=f_{i1}+f_{i2}y$ with $f_{i1}$ and $f_{i2}$ belonging to $\mathbb{F}_{q}(x)$ for $i=1,2$. For any rational place $P\neq P_{\infty}$, the matrix induced by $f_{i}$ on $\mathcal{O}^{\prime}_{P}$ over $\mathcal{O}_{P}$ is
$\begin{bmatrix}f_{i1},&f_{i2}\\ f_{i2}f_{\mathcal{E}},&f_{i1}\end{bmatrix}$.
Let $L=f_{1}+f_{2}\sigma$. Then we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}f_{11}+f_{21},&f_{12}+f_{22}\\ (f_{12}-f_{22})f_{\mathcal{E}},&f_{11}-f_{21}\end{bmatrix}$
and
$\det(\varepsilon_{P}(L))=f_{11}^{2}-f_{21}^{2}-(f_{12}^{2}-f_{22}^{2})f_{\mathcal{E}}.$
For positive integers $k$ and $k_{1}$, define sum-rank codes $C_{sr,1}$ and $C_{sr,2}$ as follows:
$C_{sr,1}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{\infty}),f_{2}\in\mathcal{L}(k_{1}Q_{\infty})\}$
and
$C_{sr,2}:=\{(\bar{\varepsilon}_{Q_{1}}(L),\ldots,\bar{\varepsilon}_{Q_{s}}(L))|L=f_{1}+f_{2}\sigma,f_{1}\in\mathcal{L}^{*}_{k_{1}}(kQ_{0,1}),f_{2}\in\mathcal{L}(k_{1}Q_{0,2})\}$
where $Q_{0,1}$ and $Q_{0,2}$ are rational places lying over $Q_{0}$. By applying the Singleton bound together with Theorems IV.1 and IV.2, we obtain
$4s-2k\leq 2d_{sr,1},2d_{sr,2}\leq 4s-k+2,$
where $s\leq q$. The following two examples are given by SageMath [44].
Example IV.1.
Let $q=7$ and consider the elliptic curve $\mathcal{E}:y^{2}=x^{3}+3$ defined over $\mathbb{F}_{q}$. The rational places of $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ are
$\{P_{0},P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$
where $P_{\alpha}$ corresponds to the prime element $x-\alpha$. For $k=6$ and $k_{1}=3$, the space $\mathcal{L}(6Q_{\infty})$ has basis $\{1,x,y,x^{2},xy,x^{3}\}$. The code $C_{sr,1}(D,6Q_{\infty})$ has parameters $[28,6,8]$. Taking $f_{1}=3$ and $f_{2}=y$, we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}3,&1\\ -(x^{3}+3),&3\end{bmatrix}$
with
$\det(\varepsilon_{P}(L))=x^{3}+5.$
Since $x^{3}+5$ is irreducible over $\mathbb{F}_{q}[x]$, the function $\det(\varepsilon_{P}(L))$ has no rational zeros. Consequently, the codeword
[…]
achieves the maximum sum-rank weight of $14=2s$. Alternatively, taking $f_{1}=1$ and $f_{2}=x^{3}$, we obtain
$\varepsilon_{P}(L)=\begin{bmatrix}x^{3}+1,&0\\ 0,&1-x^{3}\end{bmatrix}$
with
$\det(\varepsilon_{P}(L))=1-x^{6}.$
All places in $\{P_{1},P_{2},P_{3},P_{4},P_{5},P_{6}\}$ are zeros of $\det(\varepsilon_{P}(L))$. Thus, the codeword
[…]
has a sum-rank weight that attains the lower bound.
Example IV.2.
Consider the same finite field $\mathbb{F}_{q}$ and elliptic curve $\mathcal{E}$ as in the previous example. We choose the place $P_{1}$, and let $Q_{1,1}$ and $Q_{1,2}$ be the places of $\mathcal{E}/\mathbb{F}_{q}$ lying over $P_{1}$. We take the rational places
$\{P_{0},P_{2},P_{3},P_{4},P_{5},P_{6}\}.$
Let $D^{\prime}=P_{0}+P_{2}+P_{3}+P_{4}+P_{5}+P_{6}$. Then the code $C_{sr,2}(D,6P_{1})$ has parameters $[24,6,6]$ with sum-rank weight distribution:
$A_{0}=1,A_{6}=36,A_{7}=144,A_{8}=1542,A_{9}=7944,$
$A_{10}=26904,A_{11}=46959,A_{12}=34122$
where $A_{i}$ counts the number of codewords with sum-rank weight $i$.
IV-D The Gilbert-Varshamov-like Bound
In this section, we asymptotically compare our codes with the GV-like bound for sum-rank metric codes. Suppose that $\{F_{i}\}_{i\in\mathbb{N}}$ is a tower of function fields such that $F_{i+1}/F_{i}$ is a Galois extension of degree 2 for all $i\in\mathbb{N}$. Denote by $\{g_{i}\}_{i\in\mathbb{N}}$ and $\{N(F_{i})\}_{i\in\mathbb{N}}$ the genus and the number of rational places of $F_{i}$, respectively. Following the construction from previous sections, with a sequence of integers $\{k_{i}\}_{i\in\mathbb{N}}$ such that $2g_{i}-1\leq k_{i}\geq 2N(F_{i})$, we obtain a sequence of sum-rank metric codes $\{C_{i}\}_{i\in\mathbb{N}}$ with parameters
  • code length: $s_{i}\leq 4N(F_{i})$,
  • dimension: $\ell_{k_{i}}=k_{i}-g_{i}+1$,
  • minimal sum-rank distance: $d_{sr,i}\geq 2s_{i}-k_{i}$.
Therefore, we have
$d_{sr,i}+\ell_{k_{i}}\geq 2s_{i}-g_{i}+1.$
By the Hasse-Weil lower bound over $\mathbb{F}_{q}$, if either $N(F_{i})$ or $g_{i}$ tends to infinity, then so does the other.
Ihara [20] observed that the Hasse-Weil bound is no longer optimal when the genus of a function field is sufficiently large with respect to $q$. Let $A(q)$ be the Ihara’s constant, defined as
[…]
The Drinfeld-Vlǎduţ bound [47] states that $A(q)\leq\sqrt{q}-1$. If $q$ is a square, then Drinfeld and Vlǎduţ proved that equality holds. Tsfasman, Vlǎduţ, and Zink [45] provided an example of curves that attain this bound employing modular curves. Garcia and Stichtenoth [14] presented a more explicit tower of function fields. Using their technology, we can present a chain of quadratic function field extensions that attain the bound for some even $q$.
Proposition IV.1.
Let $q=2^{2^{h}}$ with $h\geq 2$, and $F_{0}=\mathbb{F}_{q}(z_{0})$ be the rational function field. Suppose that $t=2^{h-1}$ and $\{F_{i}\}_{i\in\mathbb{N}}$ is a tower of function fields such that $F_{i+1}=F_{i}(z_{i+1})$, where $z_{i}$ satisfies the equations:
  • for $j\geq 0$, $z_{jt+1}^{2}+z_{jt+1}=x_{j}^{2^{t}+1}$, where $x_{0}=z_{0}$, $x_{1}=z_{t}$ and $x_{j}=\frac{z_{jt}}{x_{j-1}}$ for $j\geq 2$,
  • for $j\geq 0$ and $2\leq i\leq t$, $z_{jt+i}^{2}+z_{jt+i}=z_{jt+i-1}$.
Then
$\lim\limits_{i\to\infty}\frac{N_{i}}{g_{i}}=\sqrt{q}-1.$
Proof.
It is sufficient to show that for any $j\geq 0$, the function field $F_{jt}$ has the form $F_{(j+1)t}=F_{jt}(z_{(j+1)t})$, where $z_{(j+1)t}$ satisfies the equation
$z_{(j+1)t}^{2^{t}}+z_{(j+1)t}=x_{j}^{2^{t}+1},$
with
$x_{0}=z_{0},\ x_{1}=z_{t},\ x_{j}=\frac{z_{jt}}{x_{j-1}}\text{ for }j\geq 2.$
Then, using [14, Corollary 3.2], we obtain
$\lim\limits_{i\to\infty}\frac{N_{i}}{g_{i}}=\lim\limits_{j\to\infty}\frac{N_{jt}}{g_{jt}}=\sqrt{q}-1.$
For $j=0$, observe that
$z_{t}^{2}+z_{t}=z_{t-1},$
$z_{t}^{4}+z_{t}^{2}=z_{t-1}^{2},$
$\cdots$
$z_{t}^{2^{t}}+z_{t}^{2^{t-1}}=z_{t-1}^{2^{t-1}}.$
Recall that $t=2^{h-1}$, which implies
$z_{t}^{2^{t}}+z_{t}$ $=(z_{t-1}^{2^{t-1}}+z_{t-1}^{2^{t-2}})+\cdots+(z_{t-1}^{2}+z_{t-1})$
$=z_{t-2}^{2^{t-2}}+z_{t-2}^{2^{t-4}}+\cdots+z_{t-2}^{4}+z_{t-2}$
$=(z_{t-2}^{2^{t-2}}+z_{t-2}^{2^{t-3}})+(z_{t-2}^{2^{t-3}}+z_{t-2}^{2^{t-4}})+\cdots+(z_{t-2}^{4}+z_{t-2}^{2})+(z_{t-2}^{2}+z_{t-2})$
$=z_{t-3}^{2^{t-3}}+z_{t-3}^{2^{t-4}}+\cdots+z_{t-3}^{2}+z_{t-3}$
$=z_{t-4}^{2^{t-4}}+z_{t-4}^{2^{t-6}}+\cdots+z_{t-4}^{4}+z_{t-4}$
$=z_{t-5}^{2^{t-5}}+z_{t-5}^{2^{t-6}}+\cdots+z_{t-5}^{2}+z_{t-5}$
$\qquad\vdots$
$=z_{1}^{2}+z_{1}$
$=x_{0}^{2^{t}+1}$
Then we have $F_{t}=F_{0}(z_{t})$ with
$z_{t}^{2^{t}}+z_{t}=x_{0}^{2^{t}+1}$
The proof can be completed with similar equations for any $j\geq 1$. ∎
The above proposition shows that there exist towers of quadratic function field extensions over $\mathbb{F}_{q}$ attaining the Drinfeld-Vlǎduţ bound for some square $q$. For any sum-rank metric code $C_{i}\in\mathbb{F}_{q}^{2\times 2}$ with $i\in\mathbb{N}$, its rate $\mathcal{R}_{i}$ and relative minimum distance $\delta_{sr,i}$ are defined as
$\mathcal{R}_{i}:=\frac{\ell_{k_{i}}}{2s_{i}},\quad\delta_{sr,i}:=\frac{d_{sr,i}}{2s_{i}}.$
Theorem IV.3.
Suppose that $q$ is a square and there exists a tower $\{F_{i}\}_{i\in\mathbb{N}}$ of quadratic function field extensions over $\mathbb{F}_{q}$ attaining the Drinfeld-Vlǎduţ bound. For any real numbers $\mathcal{R}_{sr},\delta_{sr}\in(0,1)$ satisfying
$\mathcal{R}_{sr}<1-\delta_{sr}-\frac{1}{2(\sqrt{q}-1)}.$ (1)
There exists a sum-rank metric code with rate at least $\mathcal{R}_{sr}$ and relative minimum distance at least $\delta_{sr}$.
Proof.
We only need to consider the sequence of sum-rank metric codes $\{C_{i}\}_{i\in\mathbb{N}}$ constructed from the tower $\{F_{i}\}_{i\in\mathbb{N}}$ via Theorem IV.1 or IV.2. For any $i\in\mathbb{N}$, the parameters of $C_{i}$ are
$s\leq 4N(F_{i}),\ell_{k_{i}},d_{sr,i}\geq 2s_{i}-k_{i}].$
Denote the rate $\mathcal{R}_{i}$ and relative minimum distance $\delta_{sr,i}$ as above. Choose $k_{i}$ such that $\mathcal{R}_{i}\geq\mathcal{R}_{sr}$ for all $i$ and $\lim\limits_{i\to\infty}\mathcal{R}_{i}=\mathcal{R}_{sr}$. Then we have
$\mathcal{R}_{sr}=\lim\limits_{i\to\infty}\mathcal{R}_{i}\geq 1-\lim\limits_{i\to\infty}\delta_{sr,i}-\frac{1}{2(\sqrt{q}-1)}.$
Thus for sufficiently large $i$, the code $C_{i}$ satisfies the desired parameters. ∎
Note that the GV-like bound for $n=m=2$ has the form
$\mathcal{R}_{sr}<$ $(\delta_{sr}-1)^{2}-\frac{\delta_{sr}}{2}\log_{q}(1+\frac{1}{2\delta_{sr}})-\frac{\log_{q}(1+2\delta_{sr})}{4}$
$-\frac{\log_{q}(\gamma_{q})}{4}+o(1).$ (2)
For certain values of $q$ such that there exists a tower of quadratic function field extensions attaining the Drinfeld-Vlǎduţ bound, some families of our sum-rank metric codes surpass the GV-like bound. Using the function field tower given in Proposition IV.1, we plot the comparisons between equations (1) with (IV-D) for $q=2^{4}$, $q=2^{8}$ and $q=2^{16}$ in Figs. 2, 3, and 4, respectively.
Figure 2: $q=2^{4}$
Figure 3: $q=2^{8}$
Figure 4: $q=2^{16}$
V Conclusion
In this paper, we presented two constructions of $2\times 2$ sum-rank metric codes via quadratic Galois extensions of algebraic function fields. By expressing the functions of extension fields to matrices over the base fields, we determined the parameters of these codes. Furthermore, we provided illustrative examples based on elliptic function fields.
To the best of our knowledge, these constructions are new. The only existing result on AG codes in the sum-rank metric was given by E. Berardini and X. Caruso [4]. In contrast to their work, our approach relies more heavily on the structures of function field extensions. In this paper, we established that $\det(\varepsilon_{P}(L))\neq 0$ in the special case where $f_{1}=0$. The primary challenge in extending these constructions to $n\times n$ matrices lies in guaranteeing $\det(\varepsilon_{P}(L))\neq 0$. Through SageMath computations, we verify that $L(g)$ has trivial kernel for certain special function fields and places when $n\geq 3$. Consequently, in our future work, we will investigate more general constructions.
Regarding the sum-rank distance of our AG codes in the sum-rank metric, for a fixed dimension kg(F)+1k-g(F)+1, we have
$\max d_{sr}-\min d_{sr}\approx\frac{k+g(F)+1}{2}.$
This raises an interesting problem: adapting the techniques used in constructing MDS AG codes to the sum-rank metric to obtain MSRD AG codes. Building on the approach for AG codes with narrow bounds (such as elliptic codes [11]), we identify this problem as an important objective for our subsequent research.
VI Acknowledgment
This work is supported by Natural Science Foundation of Guangdong Province (No. 2025A1515011764), the National Natural Science Foundation of China (No. 12441107) and Guangdong Provincial Key Laboratory of Information Security Technology (No. 2023B1212060026).